OpenStudy (thedj4jc):

PLEASE HELP!!!!!! Find the square root of 25 over 49.

OpenStudy (thedj4jc):

2. Simplify the square root of 36a^10b^2

OpenStudy (thedj4jc):

Answer options for the first problem 12 0.7142857 5/7 0.510204

OpenStudy (thedj4jc):

2. 6a^5b^5 6a^5b 6ab^5

OpenStudy (anonymous):

Is it \(\dfrac{\sqrt{25}}{49}\)

OpenStudy (anonymous):

Or is it \(\sqrt{\dfrac{25}{49}}\)

OpenStudy (thedj4jc):

The second one, yes!!

OpenStudy (anonymous):

Oh, just find the square roots of the numbers. Square root of 25 = ? Square root of 49 = ?

OpenStudy (thedj4jc):

5 and 7

OpenStudy (thedj4jc):

Thank you!! Do you have any idea on the second problem?

OpenStudy (anonymous):

Yes, so we have: \(\sf \dfrac{5}{7}\)

OpenStudy (thedj4jc):

Yes! Thank you!!

OpenStudy (anonymous):

Yep, find the square root of 36.

OpenStudy (thedj4jc):

6

OpenStudy (anonymous):

And we divide the exponents by 2.

OpenStudy (anonymous):

Yep. \(\sf a^{10} \rightarrow a^{10 \div 2}\)

OpenStudy (anonymous):

\(\sf b^2 \rightarrow b^{2 \div 2}\)

OpenStudy (thedj4jc):

So, that would be B. Cause 10 divided by 2 equals 5 and 2 divide by 2 equals 1.

OpenStudy (anonymous):

Yep, you got it. \(\sf 6a^5b\)